# Stairway to Success: An Online Floor-Aware Zero-Shot Object-Goal Navigation Framework via LLM-Driven Coarse-to-Fine Exploration

### v3

@@ I INTRODUCTION
Modern service and delivery robots are expected to operate in multi-floor buildings, from homes to offices. A simple command like “find the TV” becomes challenging if the target is on another floor, requiring not only object recognition but also navigating across floors and reasoning vertically, particularly in unexplored buildings. This highlights a critical gap in robotics: Object-Goal Navigation (OGN) in multi-floor settings. However, existing online navigation systems often fail in these scenarios, limiting their real-world applications [1].
To quantify the importance of multi-floor navigation, we analyze two widely-used OGN benchmarks, HM3D [2] and MP3D [3]. Over half of validation episodes involve multi-level buildings, with up to 28% requiring explicit floor transitions (see Sec. V-A). However, most OGN methods assume single-floor operation. While multi-floor SLAM systems [4, 5] improve localization, they lack semantic reasoning for goal-driven navigation. Hierarchical mapping approaches [6] operate online but prioritize mapping over efficient exploration.
Moreover, recent OGN methods fall into two categories: (i) learning-based approaches [7, 8, 9, 10], which achieve strong performance in familiar environments but require expensive data collection and retraining for unseen settings, and (ii) Zero-Shot Object-Goal Navigation (ZS-OGN) approaches [11, 12, 13, 14, 15, 16, 17, 18]. While ZS-OGN methods address the generalization limitation by operating without retraining, they still often neglect cross-floor planning and fail to handle multi-floor navigation effectively. Even recent multi-floor methods like MFNP [19] restrict exploration with one-way stair transitions, motivating our online multi-floor navigation method as shown in Fig. 1.
@@
Beyond multi-floor navigation, efficient planning poses another critical challenge for ZS-OGN. While Large Language Models (LLMs) have shown promise as high-level planners[15, 18, 17, 16], frequent API calls lead to high costs and slow execution. An alternative is to use Vision-Language Models (VLMs) for local matching, as demonstrated by VLFM [14]. However, while its greedy strategy is faster, it lacks global planning, leading to oscillation and suboptimal paths.
To address these limitations, we propose ASCENT, the first online framework that unifies dynamic multi-floor mapping with LLM-driven spatial reasoning for zero-shot navigation. Our approach presents a Multi-Floor Abstraction module that captures floor connectivity and enables inter-floor reasoning, allowing the robot to explore unseen buildings while maintaining semantic consistency across floors. By introducing a Coarse-to-Fine Reasoning module, we dramatically reduce the need for frequent LLM calls-by over 90% compared to prior planners (see Sec. V-C). This method leverages the VLM for coarse assessments while the LLM-based planner handles fine-grained decisions (e.g., floor/region selections), striking a balance between real-time performance and reasoning capability.
@@ II-A Zero-Shot Object-Goal Navigation
ZS-OGN aims to find a target object in an unknown environment without any task-specific training. Early approaches, such as ZSON [11], leveraged VLMs like CLIP [20] to transfer knowledge from Image-Goal to Object-Goal Navigation. Based on this, PSL [12] addressed semantic neglect by prioritizing goal-relevant regions. Subsequent modular methods also showed promise: ActPept [13] jointly modeled geometric and semantic cues for object association, while VLFM [14] introduced vision-language frontier maps for semantic-aware exploration. In contrast to learning-based methods [7, 8, 9, 10], ZS-OGN methods avoid expensive training. However, most existing ZS-OGN methods assumed single-floor environments and lacked explicit mechanisms for multi-floor reasoning.
II-B LLM-Based Planners for Object-Goal Navigation
Recent ZS-OGN work has turned to LLMs for high-level planning. L3MVN [15] used LLMs to generate long-horizon plans, and SG-Nav [16] constructed scene graphs for contextual memory. InstructNav [17] proposed a dynamic chain-of-navigation prompting strategy, and PixNav [18] employed LLMs as planners to guide room-level exploration. Despite their strong reasoning, these methods relied on frequent LLM API calls, causing high computational latency and limited real-time applicability. ASCENT addresses the limitations with the Coarse-to-Fine Reasoning module, which significantly reduces LLM invocation while improving planning quality.
@@ II-C Multi-Floor Navigation
Navigating across floors poses unique challenges for mobile robots, such as SLAM failures in repetitive or texture-poor environments [4, 5]. While some methods assumed access to floor plans [21] or pre-built topological graphs [6], such requirements hinder deployment in unexplored buildings. MFNP [19] was the first zero-shot approach for multi-floor OGN, but it enforces one-way floor transitions, thus preventing floor revisiting, which is a critical limitation in real-world navigation. ASCENT addresses this gap by introducing an online Multi-Floor Abstraction module that enables bidirectional exploration and cross-floor reasoning without pre-built maps.
III PROBLEM FORMULATION
We tackle the OGN task, where a robot $r$ starts from an initial position $p_{0}\in\mathbb{R}^{3}$ in an unexplored environment. The robot only has access to an egocentric RGB-D camera and an odometry sensor that provides position $p$ relative to $p_{0}$. At each timestep $t$, the robot selects discrete actions $a_{t}\in\mathcal{A}$ to generate a path $\tau$ towards discovering any instance of a target object category $\mathcal{O}$. The action space $\mathcal{A}$ consists of: MOVE FORWARD (0.25m), TURN LEFT ($30^{\circ}$), TURN RIGHT ($30^{\circ}$), LOOK UP ($30^{\circ}$), LOOK DOWN ($30^{\circ}$), and STOP. The timestep limit $T$ is $500$ steps. Following the oracle-visibility criterion of OGN evaluation [22], an episode is successful if:
IV METHODOLOGY
@@
To achieve this dual objective, ASCENT incorporates two key components. The Multi-Floor Abstraction module minimizes the exploitation cost $c_{\mathrm{goal}}$ by ensuring multi-floor accessibility and efficient goal-reaching (see Sec. IV-A). The Coarse-to-Fine Reasoning module reduces the exploration cost $c_{\mathrm{expl}}$ by prioritizing high-value frontiers (see Sec. IV-B).
IV-A Multi-Floor Abstraction
To support online navigation in multi-floor environments, our vision-based method eliminates the need for dedicated vertical sensors and abstracts the complex 3D environment into a manageable, multi-layered representation to facilitate efficient planning. This abstraction simplifies both intra-floor navigation and inter-floor transitions by converting raw RGB-D observations into structured semantic representations that enable efficient cross-floor planning. We achieve these core innovations through two key designs: BEV Mapping Representations and Multi-Floor Topology Modeling.
IV-A1 BEV Mapping Representations
Our approach builds on the concept of using Bird’s Eye View (BEV) representations for frontier-based navigation, inspired by prior work such as VLFM. Frontiers are defined as the midpoints of the boundaries between explored and unexplored areas, denoted as $\mathcal{F}$. For intra-floor exploration, we use two complementary BEV mapping representations as follows:
Exploration Value Map $\mathcal{M}_{\mathrm{val}}$: This map guides efficient local exploration by integrating a Semantic Similarity Map $\mathcal{M}_{\mathrm{ss}}$ with a proximity-based Exploration Cost Map $\mathcal{M}_{\mathrm{ec}}$. This design ensures the robot first fully exploits local, high-value areas before making high-level decisions. The total value for the $i$-th frontier $\mathcal{F}_{i}$} is defined as:
@@ IV-A1 BEV Mapping Representations
where $d_{i}$ represents the Euclidean distance from frontier $\mathcal{F}_{i}$ to the robot’s position at the current moment $t$, $d_{\theta}$ is a distance threshold, and $\mathcal{M}_{\mathrm{ss}}$ aggregates values from $[0,t]$. Intuitively, this design encourages the robot to prioritize frontiers that are both semantically relevant (high $\mathcal{M}_{\mathrm{ss}}$) and spatially accessible (small $d_{i}$), ensuring that local, promising areas are exploited before distant or cross-floor exploration.
Stair-Aware Obstacle Map $\mathcal{M}_{\mathrm{obs}}$: This is a binary occupancy grid constructed from depth data, which also aggregates values from $[0,t]$. Unlike traditional maps that treat stairs as impassable, ours dynamically re-labels stair-like structures as traversable after they are verified by the stair detection process. This re-labeling provides the topological basis for multi-floor navigation, allowing to plan paths including stair traversal.
IV-A2 Multi-Floor Topology Modeling
@@
Figure 3: Stair Detection module. ASCENT detects upward stairs (top) and infers downward stairs using depth-based analysis (bottom).
Cross-Floor Transition. When a valid stair region is identified, the robot establishes a specific frontier $\mathcal{F}_{\mathrm{stair}}$ at its midpoint. When the robot makes a cross-floor decision, it first navigates toward $\mathcal{F}_{\mathrm{stair}}$ until reaching the stair area. Then its navigation target switches to a dynamic intermediate waypoint positioned at a distance of $d_{\mathrm{stair}}=0.8\,\text{m}$ ahead of its current pose. This encourages continuous forward movement while preventing premature stopping or localization drift. The low-level locomotion leverages a pre-trained point-goal navigation (PointNav) policy without relying on global maps.
For unexplored stairs, the climbing continues until the robot exits the opposite side. For previously traversed stairs, the robot uses recorded start/end points for direct navigation. If climbing exceeds $\frac{T}{10}$ steps, the robot reverses to its previous position and marks the stair as impassable on $\mathcal{M}_{\mathrm{obs}}$.
@@
IV-B Coarse-to-Fine Reasoning
To enable context-aware exploration in multi-floor environments, we introduce a two-stage reasoning pipeline: Coarse-Grained Assessment identifies high-value frontiers efficiently, and Fine-Grained Decision performs deeper contextual reasoning to select the final target. Our method is zero-shot, requiring no explicit training on the target task. To align exploration with object distributions, we optionally leverage statistical priors from the training split, analogous to how learning-based methods implicitly capture in-domain knowledge. This guides efficient multi-floor online reasoning while preserving the zero-shot nature of our system.
IV-B1 Coarse-Grained Assessment
@@
IV-B2 Fine-Grained Decision
@@ V EXPERIMENTS
V-A Experimental Settings
Datasets. We evaluate our method on HM3D [2] and MP3D [3], the official Habitat Challenge benchmarks for OGN [23]. HM3D contains 2,000 validation episodes across 20 scenes with 6 object categories (all COCO classes), while MP3D includes 2,195 validation episodes across 11 scenes with 21 object categories (many non-COCO, open-vocabulary classes). As shown in Fig. 5, a significant portion of these environments are multi-floor: roughly 65% of validation scenarios in HM3D and 73% in MP3D. More importantly, a notable portion of episodes require cross-floor navigation, with targets located on different floors from the starting point in about 28% of episodes in HM3D and 15% in MP3D. These statistics establish the critical need for multi-floor navigation.
@@ V-A Experimental Settings
Metrics. We adopt two standard metrics from OGN evaluation [22]: Success Rate (SR) and Success weighted by Path Length (SPL) to measure navigation performance and path efficiency. All results are reported in percentages.
TABLE I: Comparisons with SOTAs on OGN task. The table contrasts supervised and zero-shot methods on HM3D and MP3D datasets across the “Success Rate” (SR) and “Success Weighted by Path Length” (SPL) metrics. We introduce columns for Vision and Language models to specify the Instruction Interpolator components used by each approach. This highlights the SOTA performance of our approach in open-vocabulary and multi-floor online navigation scenarios.
Method Venue Instruction Interpolator HM3D [2] MP3D [3]
@@
Single- Floor RIM [7] IROS’23 - - $57.8$ $27.2$ $50.3$ $17.0$
OVG-Nav [10] RAL’24 - - - - $35.8$ $12.3$
Multi- Floor PIRLNav [8] CVPR’23 - - $64.1$ $27.1$ - -
XGX [9] ICRA’24 - - $72.9$ $35.7$ - -
@@
Single- Floor ZSON [11] NeurIPS’22 CLIP [20] - $25.5$ $12.6$ $15.3$ $4.8$
L3MVN [15] IROS’23 - [Uncaptioned image] GPT-2 [24] $50.4$ $23.1$ $34.9$ $14.5$
PSL [12] ECCV’24 CLIP [20] - $42.4$ $19.2$ - -
@@
PixNav [18] ICRA’24 LLaMA-Adapter [25] [Uncaptioned image] GPT-4 [26] $37.9$ $20.5$ - -
VLFM [14] ICRA’24 BLIP-2 [27] - $52.5$ $30.4$ $36.4$ $17.5$
SG-Nav [16] NeurIPS’24 LLaVA-1.6-7B [28] [Uncaptioned image] GPT-4 [26] $54.0$ $24.9$ $40.2$ $16.0$
ActPept [13] RAL’24 GraphSAGE [29] - - - $39.8$ $17.4$
InstructNav [17] CoRL’24 - [Uncaptioned image] GPT-4V [26] $58.0$ $20.9$ - -
@@
Multi- Floor MFNP [19] ICRA’25 [Uncaptioned image] Qwen-VLChat-Int4 [30] [Uncaptioned image] Qwen2-7B [31] $58.3$ $26.7$ $41.1$ $15.4$
@@
ASCENT Ours BLIP-2 [30] [Uncaptioned image] Qwen2.5-7B [32] $\mathbf{65.4}$ $\mathbf{33.5}$ $\mathbf{44.5}$ $\mathbf{15.5}$
@@
Implementation Details. Our experiments are conducted in the Habitat simulator [33]. For object perception, we use D-FINE [34] for COCO-class objects and Grounding-DINO [35] for open-vocabulary detection, with the same confidence thresholds as VLFM for fair comparison: $\epsilon_{1}=0.8$ (COCO) and $\epsilon_{1}=0.4$ (open-vocabulary). The stair pixel proportion threshold $\epsilon_{2}$ is set to $0.25$ empirically. Object contours are segmented by Mobile-SAM [36], and stair semantic segmentation is performed using a pre-trained RedNet [37] on RGB-D inputs. For frontier reasoning, we employ BLIP-2 [27] for semantic similarity, Places365 [38] for scene classification, RAM [39] for object tagging, and Qwen2.5-7B-Instruct [32] for LLM-based reasoning. All experiments run on a single node with two NVIDIA RTX 3090 GPUs.
V-B Main Results

### v4

@@ I INTRODUCTION
Modern service and delivery robots are expected to operate in multi-floor buildings, from homes to offices. A simple command like “find the TV” becomes challenging if the target is on another floor, requiring not only object recognition but also navigating across floors and reasoning vertically, particularly in unexplored buildings. This highlights a critical gap in robotics: Object-Goal Navigation (OGN) in multi-floor settings. However, existing online navigation systems often fail in these scenarios, limiting their real-world applications.
To quantify multi-floor navigation’s importance, we analyze two widely-used OGN benchmarks, HM3D [ramakrishnan2021habitat] and MP3D [chang2017matterport3d]. Over half of validation episodes involve multi-floor buildings, with up to 28% requiring explicit floor transitions (see Sec. IV-A). However, most OGN methods assume single-floor use. While multi-floor SLAM systems [chung2024nv] improve localization, they lack semantic reasoning for OGN in unseen scenes.
Moreover, recent OGN methods fall into two categories: (i) learning-based approaches [chen2023object, ramrakhya2023pirlnav, wasserman2024exploitation, yoo2024commonsense], which achieve strong performance in familiar environments but require expensive data collection and retraining for unseen settings, and (ii) Zero-Shot Object-Goal Navigation (ZS-OGN) approaches [majumdar2022zson, guo2024object, yokoyama2024vlfm, yu2023l3mvn, yin2025sg, longinstructnav, cai2024bridging]. While ZS-OGN methods address the generalization limitation by operating without retraining, they still often neglect cross-floor planning and fail to handle multi-floor navigation effectively. Even recent multi-floor methods like MFNP [zhang2024multi] restrict exploration with one-way stair transitions, motivating our online multi-floor navigation method as shown in Fig. 1.
@@
Beyond multi-floor navigation, efficient planning poses another critical challenge for ZS-OGN. While Large Language Models (LLMs) have shown promise as high-level planners[yu2023l3mvn, cai2024bridging, longinstructnav, yin2025sg], frequent API calls lead to high costs and slow execution. An alternative is to use Vision-Language Models (VLMs) for local matching, as demonstrated by VLFM [yokoyama2024vlfm]. However, while its greedy strategy is faster, it lacks global planning, leading to oscillation and suboptimal paths.
To address these limitations, we propose ASCENT, the first online framework that unifies dynamic multi-floor mapping with LLM-driven spatial reasoning for zero-shot navigation. Our approach presents a Multi-Floor Abstraction module that captures floor connectivity and enables inter-floor reasoning, allowing the robot to explore unseen buildings while maintaining semantic consistency across floors. By introducing a Coarse-to-Fine Reasoning module, we dramatically reduce the need for frequent LLM calls-by over 90% compared to prior planners (see Sec. IV-C). This method leverages the VLM for coarse assessments while the LLM-based planner handles fine-grained decisions (e.g., floor/region selections), striking a balance between real-time performance and reasoning capability.
@@ II-A Zero-Shot Object-Goal Navigation
ZS-OGN aims to find a target object in an unknown environment without any task-specific training. Early approaches, such as ZSON [majumdar2022zson], leveraged VLMs like CLIP [radford2021learning] to transfer knowledge from Image-Goal to Object-Goal Navigation. Subsequent methods introduced modular designs: ActPept [guo2024object] jointly modeled geometric and semantic cues, VLFM [yokoyama2024vlfm] introduced vision-language frontier maps, and STRIVE [zhu2025strive] proposed structured representations with VLM-guided navigation. However, most existing methods assumed single-floor environments and lacked explicit multi-floor mechanisms.
II-B LLMs for Object-Goal Navigation
Recent ZS-OGN work has turned to LLMs for high-level planning. L3MVN [yu2023l3mvn] used LLMs to generate long-horizon plans, and SG-Nav [yin2025sg] constructed scene graphs for contextual memory. InstructNav [longinstructnav] proposed dynamic chain-of-navigation prompting, and PixNav [cai2024bridging] employed LLMs for room-level exploration. Beyond navigation planning, OpenFMNav [kuang2024openfmnav] and ApexNav [zhang2025apexnav] leverage LLMs for instruction parsing and detection enhancement respectively. Despite their strong reasoning, frequent LLM API calls cause high computational latency. ASCENT addresses this limitation through Coarse-to-Fine Reasoning, significantly reducing LLM invocations while improving planning quality.
@@ II-C Multi-Floor Navigation
Navigating across floors poses unique challenges, such as SLAM failures in repetitive or texture-poor environments [chung2024nv]. Early methods either assumed floor plans or relied on pre-built topological structures [kim2024development]. For example, Werby et al. [werby23hovsg] proposed hierarchical scene graphs for language-grounded navigation but required offline map construction. Recently, MFNP [zhang2024multi] pioneered multi-floor ZS-OGN but suffered from spatial map overlap, one-way stair constraints, and heuristic floor switching. We address these through per-floor representations, bidirectional traversal, and hierarchical planning.
III METHODOLOGY
@@
To achieve this dual objective, ASCENT incorporates two key components. The Multi-Floor Abstraction module minimizes the exploitation cost $c_{\mathrm{goal}}$ by ensuring multi-floor accessibility and efficient goal-reaching (see Sec. III-A). The Coarse-to-Fine Reasoning module reduces the exploration cost $c_{\mathrm{expl}}$ by prioritizing high-value frontiers (see Sec. III-B).
III-A Multi-Floor Abstraction
To support online navigation in multi-floor environments, our vision-based method eliminates the need for dedicated vertical sensors and abstracts complex 3D environments into a multi-layered representation to facilitate efficient planning. These core innovations through two key designs: BEV Mapping Representations and Multi-Floor Topology Modeling.
III-A1 BEV Mapping Representations
Our approach builds on the concept of using Bird’s Eye View (BEV) representations for frontier-based navigation, inspired by VLFM. Frontiers are defined as the midpoints of the boundaries between explored and unexplored areas, denoted as $\mathcal{F}$. For intra-floor exploration, we use two complementary representations:
Exploration Value Map $\mathcal{M}_{\mathrm{val}}$: This map guides efficient local exploration by integrating a Semantic Similarity Map $\mathcal{M}_{\mathrm{ss}}$ with a proximity-based Exploration Cost Map $\mathcal{M}_{\mathrm{ec}}$. Compared to VLFM, which only use $\mathcal{M}_{\mathrm{ss}}$, it ensures the robot fully exploits local, high-value areas before distant exploration. The total value for the $i$-th frontier $\mathcal{F}_{i}$} is defined as:
@@ III-A1 BEV Mapping Representations
where $d_{i}$ represents the Euclidean distance from frontier $\mathcal{F}_{i}$ to the robot’s position at the current moment $t$, $d_{\theta}$ is a distance threshold, and $\mathcal{M}_{\mathrm{ss}}$ aggregates values from $[0,t]$. The selected frontier is converted to a 2D waypoint goal and passed to the Point-Goal Navigation (PointNav) controller, which executes navigation for up to T/10 steps until the frontier is explored or the waypoint becomes unreachable. Intuitively, this design encourages the robot to prioritize frontiers that are both semantically relevant (high $\mathcal{M}_{\mathrm{ss}}$) and spatially accessible (small $d_{i}$), ensuring that promising neighbor areas are exploited fully.
Stair-Aware Obstacle Map $\mathcal{M}_{\mathrm{obs}}$: This is a binary occupancy grid constructed from depth data, which also aggregates values from $[0,t]$. Unlike VLFM, where obstacle maps treat stairs as impassable, ours re-labels stair-like structures as traversable after stair detection process. The map also maintains persistent blacklist/cache records of failed/successful transitions, providing the topological basis for multi-floor navigation.
III-A2 Multi-Floor Topology Modeling
@@
Figure 3: Stair Detection process. ASCENT detects upward stairs (top) and infers downward stairs using depth-based analysis (bottom).
@@
Cross-Floor Transition. When a valid stair region is identified, the robot establishes a specific frontier $\mathcal{F}_{\mathrm{stair}}$ at its midpoint. When the robot makes a cross-floor decision, it first navigates toward $\mathcal{F}_{\mathrm{stair}}$ until reaching the stair area. Then its navigation target switches to a dynamic intermediate waypoint positioned at a distance of $d_{\mathrm{stair}}=0.8\,\text{m}$ ahead of its current pose. This encourages continuous forward movement while preventing premature stopping or localization drift. The low-level controller leverages a pre-trained PointNav policy.
For unexplored stairs, climbing continues until the robot exits the opposite side. For previously traversed stairs, the robot uses recorded start/end points for direct navigation. If climbing exceeds $\frac{T}{10}$ steps, the robot reverses to its previous position and marks the stair as impassable on $\mathcal{M}_{\mathrm{obs}}$.
@@
Overall, we integrate multiple feasibility checks: connectivity validation through $\mathcal{M}_{\mathrm{obs}}$, passability verification via dual-modality stair detection, and access control through blacklist/cache mechanisms, ensuring robust multi-floor navigation.
III-B Coarse-to-Fine Reasoning
To enable context-aware exploration in multi-floor environments, we introduce a two-stage reasoning pipeline: Coarse-Grained Assessment identifies high-value frontiers efficiently, and Fine-Grained Decision performs deeper contextual reasoning to select the final target. We leverage statistical priors from the training split of benchmarks to guide exploration, analogous to how learning-based methods capture in-domain knowledge without task-specific training.
III-B1 Coarse-Grained Assessment
@@
III-B2 Fine-Grained Decision
@@ IV EXPERIMENTS
IV-A Experimental Settings
Datasets. We evaluate our method on HM3D [ramakrishnan2021habitat] and MP3D [chang2017matterport3d], the official Habitat Challenge benchmarks for OGN [habitatchallenge2022]. HM3D contains 2,000 validation episodes across 20 scenes with 6 object categories (all COCO classes), while MP3D includes 2,195 validation episodes across 11 scenes with 21 object categories (many non-COCO, open-vocabulary classes). As shown in Fig. 5, a significant portion of these environments are multi-floor: roughly 65% of validation scenarios in HM3D and 73% in MP3D. More importantly, a notable portion of episodes require cross-floor navigation, with targets located on different floors from the starting point in about 28% of episodes in HM3D and 15% in MP3D. These statistics establish the critical need for multi-floor navigation.
@@ IV-A Experimental Settings
Metrics. We adopt two standard metrics from OGN evaluation [batra2020objectnav]: Success Rate (SR) and Success weighted by Path Length (SPL) to measure navigation performance and path efficiency. All results are reported in percentages.
TABLE I: Comparisons with SOTA methods. The table contrasts learning-based and zero-shot methods on HM3D and MP3D datasets across the “Success Rate” (SR) and “Success Weighted by Path Length” (SPL) metrics. We introduce columns for Vision and Language models to specify the Instruction Interpolator components used by each approach. All baseline results from their original publications.
Method Venue Instruction Interpolator HM3D [ramakrishnan2021habitat] MP3D [chang2017matterport3d]
@@
Single- Floor RIM [chen2023object] IROS’23 - - $57.8$ $27.2$ $50.3$ $17.0$
OVG-Nav [yoo2024commonsense] RAL’24 - - - - $35.8$ $12.3$
Multi- Floor PIRLNav [ramrakhya2023pirlnav] CVPR’23 - - $64.1$ $27.1$ - -
XGX [wasserman2024exploitation] ICRA’24 - - $72.9$ $35.7$ - -
@@
Single- Floor ZSON [majumdar2022zson] NeurIPS’22 CLIP [radford2021learning] - $25.5$ $12.6$ $15.3$ $4.8$
L3MVN [yu2023l3mvn] IROS’23 - [Uncaptioned image] GPT-2 [liu2019roberta] $50.4$ $23.1$ $34.9$ $14.5$
@@
PixNav [cai2024bridging] ICRA’24 LLaMA-Adapter [zhang2023llama] [Uncaptioned image] GPT-4 [achiam2023gpt] $37.9$ $20.5$ - -
VLFM [yokoyama2024vlfm] ICRA’24 BLIP-2 [li2023blip2] - $52.5$ $30.4$ $36.4$ $17.5$
SG-Nav [yin2025sg] NeurIPS’24 LLaVA-1.6-7B [liu2023visual] [Uncaptioned image] GPT-4 [achiam2023gpt] $54.0$ $24.9$ $40.2$ $16.0$
OpenFMNav [kuang2024openfmnav] NAACL-F’24 [Uncaptioned image] GPT-4V [achiam2023gpt] [Uncaptioned image] GPT-4 [achiam2023gpt] $54.9$ $24.4$ $37.2$ $15.7$
ActPept [guo2024object] RAL’24 GraphSAGE [hamilton2017inductive] - - - $39.8$ $17.4$
InstructNav [longinstructnav] CoRL’24 [Uncaptioned image] GPT-4V [achiam2023gpt] [Uncaptioned image] GPT-4 [achiam2023gpt] $58.0$ $20.9$ - -
ApexNav [zhang2025apexnav] RAL’25 BLIP-2 [li2023blip2] [Uncaptioned image] DeepSeek-V3 [liu2024deepseek] $59.6$ $33.0$ $39.2$ $\mathbf{17.8}$
@@
Multi- Floor MFNP [zhang2024multi] ICRA’25 [Uncaptioned image] Qwen-VLChat-Int4 [bai2023qwen] [Uncaptioned image] Qwen2-7B [team2024qwen2] $58.3$ $26.7$ $41.1$ $15.4$
@@
ASCENT Ours BLIP-2 [li2023blip2] [Uncaptioned image] Qwen2.5-7B [yang2024qwen2] $\mathbf{65.4}$ $\mathbf{33.5}$ $\mathbf{44.5}$ 15.5{15.5}
@@
Implementation Details. Our experiments are conducted in the Habitat [savva2019habitat]. For object detection, we use D-FINE [peng2024d] for COCO-class objects and Grounding-DINO [liu2024grounding] for open-set detection, following VLFM’s confidence thresholds. Object segmentation uses Mobile-SAM [zhang2023faster], and stair semantic segmentation uses RedNet [jiang2018rednet]. For frontier reasoning, we employ BLIP-2 [li2023blip2] for semantic similarity, Places365 [zhou2017places] for scene classification, RAM [zhang2024recognize] for object tagging, and Qwen2.5-7B-Instruct [yang2024qwen2] for LLM-based reasoning. Our method leverages two offline statistical priors: (1) Floor-level priors: distribution of goal objects across floor levels from training episodes of benchmarks. (2) Area-level priors: object-room co-occurrence probabilities from HM3D statistics. All experiments run on two NVIDIA RTX 3090 GPUs.
IV-B Main Results